OpenStudy (anonymous):

One savings account begins with $1000 and earns 6% interest, compounded monthly. Another account begins at the same time with $1200 and earns 4% interest, compounded yearly. How long will it take for the two accounts to reach the same value? Give an answer in years, rounded to the nearest whole number.

OpenStudy (anonymous):

1000(1 + 0.06 / 12)^12t = 1200(1 + 0.04)^t 1.005^12t = 1.2(1.04)^t 12t log(1.005) = log(1.2) + t log(1.04) That is my steps so far and I think I need to factor out a t but not sure

OpenStudy (anonymous):

@Hero @SolomonZelman @jim_thompson5910 @tkhunny

OpenStudy (anonymous):

@Loser66

OpenStudy (tkhunny):

Have you considered subtracting [t log(1.04)] from each side? Gather the variables to one side!

OpenStudy (anonymous):

12t log(1.005) - t log(1.04) = log(1.2)

OpenStudy (anonymous):

@tkhunny, and then factor out a t?

OpenStudy (tkhunny):

That's right. Algebra I!!

OpenStudy (tkhunny):

You can change log(1.005) - log(1.04) = log(1.005/1.04) if you like.

OpenStudy (anonymous):

12t log(1.005/1.04) = log(1.2)

OpenStudy (tkhunny):

Whoops! I missed the 12. Don't do that. \(t(12\log(1.005) - \log(1.04)) = \log(1.2)\) It's time to resort to the numerical solution.

OpenStudy (anonymous):

OK, so let me write it down the correct way

OpenStudy (anonymous):

So what I have is 12t log(1.005) - log(1.004) = log(1.2) t(12 log(1.005) - log(1.004) = log(1.2)

OpenStudy (anonymous):

Then you divide both sides by log(1.005/1.04)

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

No, you didn't undo the error. \(t = \dfrac{\log(1.2)}{12\log(1.005) - \log(1.04)}\)

OpenStudy (anonymous):

I'm not sure I follow

OpenStudy (anonymous):

If I'm subtracting two logs, then I divide them

OpenStudy (anonymous):

t = 12 log(1.005)/log(1.004) = log 1.2

OpenStudy (tkhunny):

I agree that's the idea, but that is not what we have. log(a) - log(b) = log(a/b) 12log(a) - log(b) = log(a^12) - log(b) = log((a^12)/b) Quite a different animal. Change \(12\log(1.005) = \log\left(1.005^{12}\right) = \log(1.0616778)\)

OpenStudy (anonymous):

log(1.0616778) - log(1.004) = log(1.2)

OpenStudy (anonymous):

t = log(1.2) / log(1.0616788/log(1.04)

OpenStudy (tkhunny):

Where did the "t" go? \(t = \dfrac{\log(1.2)}{\log(1.0616788)-log(1.04)} = \dfrac{log(1.2)}{\log\left(\dfrac{1.0616788}{1.04}\right)}\) Pay VERY close attention to what you are dividing and what you are subtracting.

OpenStudy (anonymous):

t is approximately 8.84 years

OpenStudy (tkhunny):

Do you believe it? It takes about 9 years for 6.17% to catch up to 4% with a 20% lead? Seems okay.

OpenStudy (anonymous):

That seems alright.

OpenStudy (tkhunny):

Well, there you go. Move on to the next one.

OpenStudy (anonymous):

Thanks, I posted it.